OpenStudy (anonymous):

The technology of a school was valued at $153,500 three years ago. It is now valued at $130,570.80. What was the depreciation percentage (%) rate over those years? (A = C(1 - r)t) A. 0.525% B. 5.20% C. 5.25% D. 5.30%

OpenStudy (aum):

\(\Large A = C(1-r)^t \) A = $130,570.80 C = $153,500 t = 3 years Substitute and solve for r.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

May have another for you :)

OpenStudy (aum):

You are welcome. I can try one more.

OpenStudy (anonymous):

what is the common ratio of this geometric sequence? (following are in 2 column chart) 1 - 27 2 - 18 3 - 12 4 - 8 5 - 16/3 Answers possibilities: -9 2/3 3/2 9

OpenStudy (aum):

For the first one, when you solve for r it will be in decimal. Multiply it by 100 to convert to percentage.

OpenStudy (anonymous):

got it. any advice on the other question? thanks!

OpenStudy (aum):

To find the common ratio, take any term and divide it by the previous term. Example, second term / first term = 18/27 = 2/3 third term / second term = 12/18 = 2/3

OpenStudy (anonymous):

thanks!